# Precession and higher order multipoles in binary black holes (and lack thereof)

### v1

@@ Abstract
The latest binary black hole population estimates argue for a subpopulation of unequal component mass binaries with spins that are likely small but isotropically distributed. This implies a non-zero probability of detecting spin-induced orbital precession and higher order multipoles moments in the observed gravitational-wave signals. In this work we directly calculate the probability for precession and higher order multipoles in each significant gravitational-wave candidate observed by the LIGO–Virgo–KAGRA collaborations (LVK). We find that only one event shows substantial evidence for precession: GW200129_065458, and two events show substantial evidence for higher order multipoles: GW190412 and GW190814; any evidence for precession and higher order multipoles in other gravitational-wave signals is consistent with random fluctuations caused by noise. We then compare our observations with expectations from population models, and confirm that current population estimates from the LVK accurately predict the number of observed events with significant evidence for precession and higher order multipole moments. In particular, we find that this population model predicts that a binary with significant evidence for precession will occur once in every $\sim 50$ detections, and a binary with significant evidence for higher order multipole moments will occur once in every $\sim 70$ observations. However, we emphasise that since substantial evidence for precession and higher order multipole moments have only been observed in three events, any population model that includes a subpopulation of binaries yielding $\sim 2\%$ of events with detectable precession and higher order multipole moments will likely be consistent with the data.
@@ I Introduction
A GW is typically decomposed into multipole moments using the $-2$ spin-weighted spherical harmonics [49]. Binary black holes with spins aligned with the orbital angular momentum, hereafter referred to as non-precessing binaries, predominantly emit radiation in the dominant $(\ell,|m|)=(2,2)$ quadrupole. However, additional power may also be radiated in higher order multipole moments, including the $(\ell,|m|)=(3,3)$ and $(4,4)$ multipoles. Since the amplitude of each higher order term is significantly smaller than the dominant quadrupole, it is challenging to clearly identify additional power within the GW strain data. However, as the binary masses become more asymmetric, the relative amplitude of the higher order multipoles increases [50]; indeed many of the higher order multipole moments amplitudes vanish for equal mass binaries, particularly the $(\ell,|m|)=(3,3)$ multipole which is otherwise the most significant for sources observed with current ground-based GW detectors [50]. Since GW190412 [38] and GW190814 [51] were the first events to have unequivocally unequal masses, they were the first signals that exhibited strong evidence for higher order multipole moments [38, 51, 52, 53, 54]. In addition marginal evidence for higher order multipoles was also reported in GW151226 [55], GW170729 [1, 56, 57], GW190519_153544 [54] and GW190929_012149 [54]. Despite strong degeneracies making it challenging to infer the mass ratio of the binary, strong evidence for the presence of the $(\ell,m)=(3,3)$ multipole was found in the population of binary black holes from the first half of the third GW observing run [54], while only weak preference was found for higher order multipoles in the first gravitational-wave catalogue (GWTC-1) [58].
@@
In this paper we take advantage of the harmonic decomposition of a GW [70] to directly calculate the probability for precession and higher order multipoles in all significant GWs observed by the LVK. We show that only one event shows substantial evidence for precession: GW200129, and two events show substantial evidence for higher order multipoles: GW190412 and GW190814. We therefore argue that any apparent evidence for precession and higher order multipoles in other GW signals is consistent with random fluctuations caused by noise at the population level; for example, in a population of 100 events, we expect one GW signal to have a 1 in 100 random fluctuation from noise alone.
We then compare our observations with expectations from population models, and show that the population model from the LVK [15] accurately predicts the number of observed events with significant evidence for precession and higher order multipole moments; we show that a binary with significant evidence for precession will occur once in every $\sim 50$ detections, and a binary with significant evidence for higher order multipole moments will occur once in every $\sim 70$ observations (when excluding GW190814 to remain consistent with current population estimates [15]), when assuming detector sensitivities during the third gravitational-wave observing run. However, we argue that since substantial evidence for precession and higher order multipole moments has only been observed in three events to date, any population model that includes a subpopulation of binaries with $\sim 2\%$ chance of detectable precession and higher order multipole moments will likely be consistent with the data.
This paper is structured as follows: in Section II we briefly review the process of identifying whether an observed GW has significant evidence for precession and higher order multipoles and in Section III.1 we identify how many GW observations have significant evidence for precession and higher order multipoles. We then investigate a correlation between the precision at which the binary’s precession phase can be measured and its observed evidence for precession in Section III.2. In Section III.3 we then calculate the probability of detecting a GW with significant evidence for precession and higher order multipole moments based on current population estimates. We finally conclude in Section IV with a summary and discussion of future directions.
@@ II Methods
As the binary black hole inspirals due to the emission of GWs, it may undergo spin-induced orbital precession leaving visible modulations in the observed GW [37]. A binary black hole undergoes precession when the spin of the binary $\mathbf{S}=\mathbf{S}_{1}+\mathbf{S}_{2}$ is misaligned with the orbital angular momentum, $\mathbf{S}\times\mathbf{L}\neq 0$. Although the strength of precession is typically characterised by the tilt angle of the binary’s orbit, $\beta$, defined as the polar angle between $\mathbf{L}$ and the total angular momentum $\mathbf{J}=\mathbf{L}+\mathbf{S}$, spin-induced orbital precession is not always observable in the detected GW signal. This is because the observability of precession in a GW signal depends on the binary’s extrinsic parameters [40, 41, 42]. For example, precession in a binary observed close to face-on, $\theta_{JN}\sim 0$, is significantly harder to detect than precession in a binary observed edge-on, $\theta_{JN}=\pi/2$ [37, 71, 72, 73, 74, 41]; precession is difficult to detect in binaries observed exactly face-on, $\theta_{JN}=0$, since the leading precession harmonics vanish [40].
According to general relativity, a GW produced by a binary black hole is fully described by two polarizations: the plus and cross polarization. These polarizations are typically decomposed into multipole moments using the $-2$ spin-weighted spherical harmonics [49]. A more significant fraction of power is radiated in multipole moments beyond the dominant quadrupole for binaries with unequal component masses [75, 76, 50]. However, as with precession, higher order multipole moments may not always be detectable in the observed GW signal since their observability depends on the binary’s extrinsic parameters: higher order multipole moments are more easily detectable in binaries observed edge-on than binaries observed close to face-on [50].
@@ II.1 The harmonic decomposition of a GW signal
Given that the observability of precession and higher order multipole moments depends on the binary’s intrinsic and extrinsic parameters, Ref. [70] most recently introduced the harmonic decomposition of a GW signal. This decomposition specifically isolates the precessing and higher order multipole contributions to the observed GW signal and allows for their amplitudes to be calculated.
The harmonic decomposition of a GW expresses the signal as an infinite sum of waveform components. However, by rewriting the sum as a power series, we understand that only a small number of these components make a significant contribution to the observed GW. As a result, by identifying the most significant waveform components, the observed GW can be greatly simplified with little loss of accuracy.
The harmonic decomposition allows us to approximate a GW signal as a sum of orthogonal waveform components,
$h\approx h^{\prime}+h_{\mathrm{p}}+\sum_{\ell,m}h_{\ell m}$ (1)
where $h^{\prime}$ is the leading order quadrupole component from an equivalent binary but with spins aligned with the orbital angular momentum, $h_{\mathrm{p}}$ is the precessing component and $h_{\ell m}$ corresponds to each higher order multipole contribution. Through the harmonic decomposition, we naturally see the effect of precession and higher order multipoles through $h_{\mathrm{p}}$ and $h_{\ell m}$. This allows us to make several key observations:
@@
    If $h_{\mathrm{p}}$ and $h_{\ell m}$ are both small in amplitude, the GW signal will be similar to the signal produced from a non-precessing binary with no observable higher order multipoles, $h\approx h^{\prime}$.
@@
    If $h_{\mathrm{p}}$ is small and $h_{\ell m}$ large, the observed GW will be similar to a non-precessing binary with observable higher order multipoles, $h\approx h^{\prime}+h_{\ell m}$.
@@
    If $h_{\mathrm{p}}$ is large and $h_{\ell m}$ small, the observed GW will be similar to the signal produced from a precessing binary with no observable higher order multipoles, $h\approx h^{\prime}+h_{\mathrm{p}}$. This will typically occur for a binary with close to equal component masses and high in-plane spins.
@@
    If $h_{\mathrm{p}}$ and $h_{\ell m}$ are both large, the observed GW will be similar to a signal produced from a precessing binary with observable higher order multipoles, $h\approx h^{\prime}+h_{\mathrm{p}}+h_{\ell m}$.
For this reason, it is useful to quantify the power, or signal-to-noise ratio (SNR), in $h_{\mathrm{p}}$ and $h_{\ell m}$, $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ respectively.
Of course, GWs are masked by noise from the GW detectors. This means that $h_{\mathrm{p}}$ and $h_{\ell m}$ may be mimicked by non-astrophysical sources, giving rise to non-zero observed SNRs. Recently, Ref. [54] laid out an algorithm for identifying the probability that the observed $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ are caused by Gaussian noise, and we refer the reader to Ref. [54] for a detailed explanation.
@@
In this work, we wish to calculate the probability that the SNR in precession, $\rho_{\mathrm{p}}$, and the SNR in higher order multipole moments, $\rho_{\ell m}$, for each observed GW signal is caused by noise. This will allow us to identify which GW signals show significant evidence for precession and higher order multipole moments, and which are simply consistent with random fluctuations caused by Gaussian noise. We will then be able to make statements about the astrophysical implications of the individual sources, and of the population as a whole.
@@
As described in Ref.[54], the probability that the observed SNR in precession, $\rho_{\mathrm{p}}$, and the SNR in higher order multipole moments, $\rho_{\ell m}$, is caused by noise can be calculated by comparing the inferred SNRs to the expected distributions from noise. Rather than measuring $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ directly from the GW strain data, as is done in Ref.[70], we can infer them from the posterior samples obtained through Bayesian inference analyses [42, 50, 54]. This can be achieved by taking advantage of the PESummary [77] and simple-pe [70] libraries. In this work we use the publicly released posterior samples, made available in the GWTC-3 data release [4], and calculate $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ by using the posterior samples obtained with the IMRPhenomXPHM [78] waveform model. GW200129 is the only signal where the IMRPhenomX waveform family is not used since it was shown that the NRSur7dq4 [79] model more accurately describes the observed GW signal [43].
@@
As described in Ref.[54], additional Bayesian inference analyses are required in order to calculate the expected $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ distributions for signals where Gaussian noise is solely responsible for producing the observed precession and higher order multipole signatures; the expected $\rho_{\mathrm{p}}$, and $\rho_{\ell m}$ distributions from Gaussian noise requires posterior samples obtained with waveform models that restricts attention to spins aligned with the orbital angular momentum, and waveform models that include only the $\ell=2$ multipoles respectively. These additional analyses are used to calculate a more informative prior for the expected distribution from noise, which accounts for the much larger volume of parameter space consistent with low SNRs in $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$; for instance an observation of an equal mass ratio binary reduces the expectation of observing higher order multipoles and precession from noise. Uninformative priors could be used for all events, but it was previously shown in Refs. [42, 50, 54] that more accurate results are obtained when using the informed prior. We obtain these additional posterior samples by analysing each observed GW signal with the IMRPhenomXHM [80] and IMRPhenomXP [78] waveform models respectively. All settings were chosen to match the original LVK analyses. For the case of GW200129 we use the same settings as those described in Ref. [43], and we obtain these additional posterior samples by analysing GW200129 with the aligned-spin limit of the NRSur7dq4, and NRSur7dq4 restricted to use just the $\ell=2$ multipoles respectively.
@@
Throughout this work we restrict attention to the observation of only the $(3,3)$ higher order multipole moment. This is because Refs. [50, 54] most recently highlighted that only the $(3,3)$ multipole is expected to contribute significantly to GW signals observed with current detector sensitivities. Consequently, throughout this work, we use a simplified version of Equation 1: $h\approx h^{\prime}+h_{\mathrm{p}}+h_{33}$. We note, however, that for binaries that are strongly precessing, or emit significant power in higher order multipoles beyond the $(\ell,|m|)=(3,3)$ harmonic, additional terms are needed in this harmonic decomposition [50, 81, 70]. Although this approximation has been rigorously tested in other works [e.g. 70], we further show the validity of this approximation for a binary with significant precession and higher order multipole contributions in Appendix A. We show that even for an extreme binary configuration, which is unlikely to be observed according to current population estimates, we see good agreement between the approximate and true GW signals.

### v2

# Rarity of precession and higher-order multipoles in gravitational-waves from merging binary black holes

@@ Abstract
The latest binary black hole population estimates argue for a subpopulation of unequal component mass binaries with spins that are likely small but isotropically distributed. This implies a non-zero probability of detecting spin-induced orbital precession and higher order multipole moments in the observed gravitational-wave signals. In this work we directly calculate the probability for precession and higher-order multipoles in each significant gravitational-wave candidate observed by the LIGO–Virgo–KAGRA collaborations (LVK). We find that only one event shows substantial evidence for precession: GW200129_065458, and two events show substantial evidence for higher-order multipoles: GW190412 and GW190814; any evidence for precession and higher-order multipole moments in other gravitational-wave signals is consistent with random fluctuations caused by noise. We then compare our observations with expectations from population models, and confirm that current population estimates from the LVK accurately predict the number of observed events with significant evidence for precession and higher-order multipoles. In particular, we find that this population model predicts that a binary with significant evidence for precession will occur once in every $\sim 50$ detections, and a binary with significant evidence for higher-order multipoles will occur once in every $\sim 70$ observations. However, we emphasise that since substantial evidence for precession and higher-order multipoles have only been observed in three events, any population model that includes a subpopulation of binaries yielding $\sim 2\%$ of events with detectable precession and higher-order multipole moments will likely be consistent with the data.
@@ I Introduction
A GW is typically decomposed into multipole moments using the $-2$ spin-weighted spherical harmonics [49]. Binary black holes with spins aligned with the orbital angular momentum, hereafter referred to as non-precessing binaries, predominantly emit radiation in the dominant $(\ell,|m|)=(2,2)$ quadrupole. However, additional power may also be radiated in higher order multipole moments (HoMs), including the $(\ell,|m|)=(3,3)$ and $(4,4)$ multipoles. Since the amplitude of each higher order term is significantly smaller than the dominant quadrupole mode, it is challenging to clearly identify additional power within the GW strain data. However, as the binary masses become more asymmetric, the relative amplitude of the HoMs increases [50]; indeed many of the HoMs amplitudes vanish for equal mass binaries, particularly the $(\ell,|m|)=(3,3)$ multipole which is otherwise the most significant for sources observed with current ground-based GW detectors [50]. Since GW190412 [38] and GW190814 [51] were the first events to have unequivocally unequal masses, they were the first signals that exhibited strong evidence for HoMs [38, 51, 52, 53, 54]. In addition marginal evidence for HoMs was also reported in GW151226 [55], GW170729 [1, 56, 57], GW190519_153544 [54] and GW190929_012149 [54]. Despite strong degeneracies making it challenging to infer the mass ratio of the binary, strong evidence for the presence of the $(\ell,m)=(3,3)$ multipole was found in the population of binary black holes from the first half of the third GW observing run [54], while only weak preference was found for HoMs in the first gravitational-wave catalogue (GWTC-1) [58].
@@
In this paper we take advantage of the harmonic decomposition of a GW [70] to directly calculate the probability for precession and HoMs in all significant GWs observed by the LVK. We show that only one event shows substantial evidence for precession: GW200129, and two events show substantial evidence for HoMs: GW190412 and GW190814. We therefore argue that any apparent evidence for precession and HoMs in other GW signals is consistent with random fluctuations caused by noise at the population level; for example, in a population of 100 events, we expect one GW signal to have a 1 in 100 random fluctuation from noise alone.
We then compare our observations with expectations from population models, and show that the population model from the LVK [15] accurately predicts the number of observed events with significant evidence for precession and HoMs; we show that a binary with significant evidence for precession will occur once in every $\sim 50$ detections, and a binary with significant evidence for HoMs will occur once in every $\sim 70$ observations (when excluding GW190814 to remain consistent with current population estimates [15]), when assuming detector sensitivities during the third gravitational-wave observing run. However, we argue that since substantial evidence for precession and HoMs has only been observed in three events to date, any population model that includes a subpopulation of binaries with $\sim 2\%$ chance of detectable precession and HoMs will likely be consistent with the data.
This paper is structured as follows: in Section II we briefly review the process of identifying whether an observed GW has significant evidence for precession and HoMs and in Section III.1 we identify how many GW observations have significant evidence for precession and HoMs. We then investigate a correlation between the precision at which the binary’s precession phase can be measured and its observed evidence for precession in Section III.2. In Section III.3 we then calculate the probability of detecting a GW with significant evidence for precession and HoMs based on current population estimates. We finally conclude in Section IV with a summary and discussion of future directions.
@@ II Methods
As the binary black hole inspirals due to the emission of GWs, it may undergo spin-induced orbital precession leaving visible modulations in the observed GW [37]. A binary black hole undergoes precession when the spin of the binary $\mathbf{S}=\mathbf{S}_{1}+\mathbf{S}_{2}$ is misaligned with the orbital angular momentum, $\mathbf{S}\times\mathbf{L}\neq 0$. In this case, both the spins and the orbital angular momentum precess around the direction of the total angular momentum. The strength of precession is typically characterised by the tilt angle of the binary’s orbit, $\beta$, defined as the polar angle between $\mathbf{L}$ and the total angular momentum $\mathbf{J}=\mathbf{L}+\mathbf{S}$, or by the scalar quantity $\chi_{\mathrm{p}}$, defined in Ref. [73], which parametrises the rate of precession through the spin components misaligned with the orbital angular momentum. However, these quantities are not sufficient to unambiguously identify if the source is precessing through Bayesian methods. For example, non-zero values of $\chi_{p}$ are often reported, even when there is no observable precession, due to the choices of priors on the spins. The visible modulations in the observed GW depend strongly on the binary’s extrinsic parameters [40, 41, 42]. For instance, precession in a binary observed close to face-on, $\theta_{JN}\sim 0$, is significantly harder to detect than precession in a binary observed edge-on, $\theta_{JN}=\pi/2$ [37, 74, 75, 76, 77, 41] since the leading precession contributions to the waveform vanish for face-on signals [40].
@@ II.1 The harmonic decomposition of a GW signal
Given that the observability of precession and HoMs depends on the binary’s intrinsic and extrinsic parameters, Ref. [70] most recently introduced the harmonic decomposition of a GW signal. This decomposition specifically isolates the leading-order contribution to the waveform from precession and HoMs and allows for their amplitudes to be calculated. Here, we briefly describe the decomposition, and refer interested readers to [78] and Appendix A of Ref. [70] for more details.
As is well known, a gravitational-wave signal can be decomposed into a series of (spin-weighted) spherical harmonics. For a spin-aligned system, it is natural to use a co-ordinate frame aligned with the orbital angular momentum when performing the decomposition. For a precessing system, the emission is also most simply described in a frame aligned with the orbital angular momentum. However, due to precession, this is not a fixed frame, but rather one that co-precesses with the binary. It has been shown that the signal from a precessing binary is well approximated by “twisting up” [79, 80, 78] the multipoles of the non-precessing counterpart, based on the evolution of $\mathbf{L}$ around $\mathbf{J}$. In effect, this twisting up procedure leads to each $(\ell,|m|)$ multipole in the co-precessing frame being split across all multipoles of the same $\ell$ in the $\mathbf{J}$-aligned frame [78]; for example the dominant $(\ell,|m|)=(2,2)$ multipole in the co-precessing frame is split across all five $\ell=2$ multipoles. By re-expressing the waveform in terms of harmonics of specific frequency, it can be shown that each multipole can be written as sum of $12\ell+$ 1 precessing harmonics, whose frequencies differ by multiples of the precession frequency. Individual harmonics do not exhibit precession, but the well known amplitude and phase modulations are obtained through neighbouring harmonics moving in and out of phase. Furthermore, the amplitudes of the precessing harmonics form a power series in $b=\tan{\beta/2}$, and for the majority of the parameter space $b\ll 1$. In [40], it was demonstrated that two harmonics are sufficient to describe the $(\ell,|m|)=(2,2)$ waveform, i.e. $h_{22}\approx h^{\prime}+h_{\mathrm{p}}$. The amplitudes of other $(\ell,|m|)$ multipoles are smaller than the $(2,2)$ and for these the leading precession harmonic suffices [70].
As a result, a GW signal in the $\mathbf{J}$-aligned frame can be expressed as,
$h\approx h^{\prime}+h_{\mathrm{p}}+\sum_{(\ell,|m|)\,\neq\,(2,2)}h_{\ell m}$ (1)
where $h_{\ell m}$ corresponds to the leading precessing harmonic for each HoM $(\ell,|m|)\neq(2,2)$. Through the harmonic decomposition, we naturally see the effect of precession and HoMs through $h_{\mathrm{p}}$ and $h_{\ell m}$. This allows us to make several key observations:
@@
    If $h_{\mathrm{p}}$ and $h_{\ell m}$ are both small in amplitude, the GW signal will be similar to the signal produced from a non-precessing binary with no observable HoMs, $h\approx h^{\prime}$.
@@
    If $h_{\mathrm{p}}$ is small and $h_{\ell m}$ large, the observed GW will be similar to a non-precessing binary with observable HoMs, $h\approx h^{\prime}+h_{\ell m}$.
@@
    If $h_{\mathrm{p}}$ is large and $h_{\ell m}$ small, the observed GW will be similar to the signal produced from a precessing binary with no observable HoMs, $h\approx h^{\prime}+h_{\mathrm{p}}$. This will typically occur for a binary with close to equal component masses and high in-plane spins.
@@
    If $h_{\mathrm{p}}$ and $h_{\ell m}$ are both large, the observed GW will be similar to a signal produced from a precessing binary with observable HoMs, $h\approx h^{\prime}+h_{\mathrm{p}}+h_{\ell m}$.
For this reason, it is useful to quantify the power, or signal-to-noise ratio (SNR), in $h_{\mathrm{p}}$ and $h_{\ell m}$, $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ respectively. Intuitively, $\rho_{\mathrm{p}}$ parameterises the amount of observable precession in a GW signal, and is what we actually measure in the data (thereby avoiding complications due to non-trivial priors). Previous studies [42] have shown that $\rho_{\mathrm{p}}$ indeed correlates strongly with the measurability of precession across the parameter space, and can be used to identify when the inference on $\chi_{\mathrm{p}}$ and $\beta$ may sensitively depend on the chosen prior: when $\rho_{\mathrm{p}}$ is small, the GW signal will be similar to the signal produced from a non-precessing binary. Similarly, $\rho_{\ell m}$ quantifies the amount of observable HoM content in the signal. Given that $\rho_{\ell m}$ increases for more asymmetric binaries observed edge-on [50], when $\rho_{\ell m}$ is small, the binary must have (close-to) equal mass components or be observed approximately face-on. This information is often difficult to extract from marginalized posterior distributions alone due to parameter degeneracies, see e.g. [30, 81].
@@
We wish to calculate the probability that the SNR in precession, $\rho_{\mathrm{p}}$, and the SNR in HoMs, $\rho_{\ell m}$, for each observed GW signal is caused by noise. This allows us to identify which GW signals show significant evidence for precession and HoMs, and which are simply consistent with random fluctuations caused by Gaussian noise. We will then be able to make statements about the astrophysical implications for the individual sources, and the population as a whole.
@@
Rather than measuring $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ directly from the GW strain data, as is done in Ref.[70], here we infer them from the posterior parameter estimates obtained through Bayesian inference analyses [42, 50, 54]. This can be achieved by taking advantage of the PESummary [82] and simple-pe [70] libraries. In this work we use the publicly released posterior samples, made available in the GWTC-3 data release [4], and calculate $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ by using the posterior samples obtained with the IMRPhenomXPHM [83] waveform model. GW200129 is the only signal where the IMRPhenomX waveform family is not used since it was shown that the NRSur7dq4 [84] model more accurately describes the observed GW signal [43].
As described in Ref.[54], the expected $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ distributions are dependent on the matched filter SNRs, $\rho^{MF}(d)$ (we hereafter drop the explicit dependence on the data $d$), of the subdominant contributions to the waveform: $h_{\mathrm{p}}$ for precession and $h_{\ell m}$ for higher order multipoles. When marginalizing the likelihood over the unknown phase, the expected posterior distribution can be shown to be,
$p(\rho|d)\propto p(\rho)\,I_{0}(\rho^{\mathrm{MF}}\rho)\exp{\left(-\frac{[\rho^{\mathrm{MF}}]^{2}+\rho^{2}}{2}\right)}\,.$ (2)
where $\rho$ denotes either the precession or HoM SNR, $p(\rho)$ is the prior distribution for $\rho$, $p(\rho|d)$ is the posterior distribution given data $d$, $I_{0}$ is the Bessel function of the first kind and $\rho^{\mathrm{MF}}$ is the matched filter SNR. The parameter estimation results provide us with the posterior distribution $p(\rho|d)$ and from this, we wish to estimate the matched filter SNR, $\rho^{MF}$. We can then calculate the probability of obtaining $\rho^{MF}$ due to noise alone, as this follows a known distribution.
@@
To calculate the expected SNR distribution for a given value of $\rho^{MF}$, the prior distribution $p(\rho)$ is needed. As described in Ref. [54], an informed prior can be used to reduce systematic biases in the distribution. The informed prior accounts for the volume of parameter space consistent with different SNRs in $\rho_{\mathrm{p}}$ and $\rho_{\ell m}$ — for instance observation of an equal mass binary reduces the expectation of observing HoMs and precession from noise, while a binary with unequal masses is much more likely to have observable power in precession or HoMs. The construction of informed priors for $\rho_{p}$ and $\rho_{\ell m}$ uses results of parameter estimates that neglect precession and higher order modes respectively. We obtain these additional posterior samples by analysing each observed GW signal with the IMRPhenomXHM [85] and IMRPhenomXP [83] waveform models. All settings were chosen to match the original LVK analyses. For the case of GW200129 we use the same settings as those described in Ref. [43], and we obtain these additional posterior samples by analysing GW200129 with the aligned-spin limit of the NRSur7dq4, and NRSur7dq4 restricted to use just the $\ell=2$ multipoles respectively.
Once the informed prior distribution has been constructed, we infer the values of $\rho_{\mathrm{p}}^{MF}$ and $\rho_{\ell m}^{MF}$ which are consistent with the observed distributions. In practice, we obtain the value of $\rho^{\mathrm{MF}}$ by minimizing the Jensen-Shannon divergence [86] between the observed $p(\rho|d)$ and the inferred distribution given a specific value of $\rho^{MF}$. Finally, we assign a p-value by identifying the probability of obtaining the required value of $\rho^{\mathrm{MF}}$ from a chi distribution with 2 degrees of freedom [54].
@@
Throughout this work we restrict attention to the observation of only the $(3,3)$ HoM. This is because Refs. [50, 54] most recently highlighted that only the $(3,3)$ multipole is expected to contribute significantly to GW signals observed with current detector sensitivities. Consequently, throughout this work, we use a simplified version of Equation 1: $h\approx h^{\prime}+h_{\mathrm{p}}+h_{33}$. We note, however, that for binaries that are strongly precessing, or emit significant power in HoMs beyond the $(\ell,|m|)=(3,3)$ harmonic, additional terms are needed in this harmonic decomposition [50, 87, 70]. Although this approximation has been rigorously tested in other works [e.g. 70], we further show the validity of the approximation for a binary with significant precession and HoM contributions in Appendix A. We show that even for an extreme binary configuration, which is unlikely to be observed according to current population estimates, we see good agreement between the approximate and full GW signals.